The enthalpy of vaporization of compound 24.2 KJ/mol At 15.0°C it has a vapor pressure of 160 mmHg , what is its vapor pressure at 25°C


1
Expert's answer
2020-07-20T05:09:23-0400

The Clausius-Clapeyron relation can be used to estimate the vapor pressure at 25°C:

"\text{ln}(\frac{p_1}{p_2}) = \frac{\Delta H_{vap}}{R}(\frac{1}{T_2}-\frac{1}{T_1})" ,

where "\Delta H_{vap}" is the enthalpy of vaporization, "T" and "p" are the temperature and the pressure, respectively and "R" is the gas constant, 8.314 J mol-1 K-1.

Making use of the equation above, we can derive the expression for "p_2" :

"p_2 = p_1\text{exp}( -\frac{\Delta H_{vap}}{R}(\frac{1}{T_2}-\frac{1}{T_1}))"

"p_2 = 160·\text{exp}( -\frac{24.2·10^3}{8.314}(\frac{1}{25+273.15}-\frac{1}{15+273.15}))"

"p_2 =225" mmHg.

Don't forget to switch from celsius to kelvin (adding 273.15) and also to switch from kJ to J for the enthalpy of vaporization (multiplying by a factor of 103).

Answer: if the enthalpy of vaporization of compound 24.2 kJ/mol and at 15.0°C it has a vapor pressure of 160 mmHg , its vapor pressure at 25°C is 225 mmHg.